OpenStudy (madeleine2):

Given the polynomial 21x 4 + 3y - 6x 2 + 34 What polynomial must be subtracted from it to obtain 29x 2 - 7 ? What polynomial must be added to it to obtain a first degree polynomial?

OpenStudy (wcrmelissa2001):

what is 21x4 is it 21x to the power of 4?

OpenStudy (wcrmelissa2001):

assuming it is.... for the first question, we apply the basic rule of subtraction: a-b=c a-c=b So for example 9-4=5 and 9-5=4 So 21x 4 + 3y - 6x 2 + 34 minus the polynomial we want to find = 29x2-7 likewise, 21x 4 + 3y - 6x 2 + 34 minus 29x2-7=the polynomial we want to find. so simply subtract 29x2-7 from 21x 4 + 3y - 6x 2 + 34 and you should get the answer